POLARIZATION OF THE SiO J=2→1 EMISSION IN ORION IRc2

\underline{RAMPRASAD RAO}, Department of Astronomy, University of Illinois, 1002 W. Green St., Urbana, IL 61801; RICHARD L. PLAMBECK and MELVYN C. H. WRIGHT, Radio Astronomy Laboratory, University of California, Berkeley, CA 94720.

We mapped the polarization structure of SiO J=2→1 emission in Orion IRc2 in both the ground state (v=0) and the vibrationally excited state (v=1). The data were obtained using the Berkeley-Illinois-Maryland-Association (BIMA) array at a frequency of 86 GHz. The angular resolution that was achieved was 0.4′′. Both the ground state and the vibrationally exited state are associated with maser emission from a disk around a young stellar object (YSO) in IRc2^{a}. The v=1 emission has 3% linear polarization which rotates across the spectral line. The fractional linear polarization and the position angle of the plane of polarization vary with time. The v=0 polarized emission seems to arise from a highly compact source but the data show that the polarization angle is constant across the spectrum. In addition, there is no time variability. The data are consistent with a rotating, expanding disk between 40 and 80 AU from a young stellar object (YSO) associated with a radio source (I). The magnetic field is entrained in the disk’s motion. At larger radii from source I, the SiO J=2→1 v=0 traces a flared disk structure on a scale of 1000 AU. The magnetic field on these scales is not strongly influenced by the star formation, or the outflow which is centered on source I.

This work was partially funded by NSF grants AST 96-13998, AST96-13999 and by the University of Illinois and the University of California.

aM. C. H. Wright, R. L. Plambeck, L. G. Mundy and L. W. Looney Astrophys. J. \underline{455}, L185 (1995).